OpenStudy (anonymous):

simplify 9(x + 3) + 2(x + 1)

mathslover (mathslover):

use distributive property to solve : a(b+c) = ab + ac 9(x+3) = 9x + 9(3) = 9x + 27 2(x+1) = 2x + 2(1) = 2x + 2 So you get : 9(x+3) + 2(x+1) = 9x + 27 + 2x + 2

mathslover (mathslover):

Combine like terms and simplify